# DressCode: Autoregressively Sewing and Generating Garments from Text Guidance

### v1

@@ 3.3 Implementation Details
We utilize the extensive sewing pattern garment dataset from [21], notable for its comprehensive range of sewing patterns and styles of garments, including shirts, hoods, jackets, dresses, pants, skirts, jumpsuits, vests, etc. Our experiments use approximately 19264 samples across 11 fundamental categories. Each garment in the dataset contains a sewing pattern file, a 3D garment mesh draped on a T-pose human model, and a rendered image. We employ the GPT-4V [1] to generate captions for garments from the rendered images of the front view and the back view. For each garment, we first prompt GPT-4V to generate its common name (e.g., hood, T-shirt, blouse) if available, followed by a request for specific geometric features (e.g., long sleeves, wide garment, deep collar), as demonstrated in Figure 3. We combine these two descriptions to form the caption for each garment. We use about 90% of the data from each category for training and the remaining for validation.
@@
We set $K=14$, and the maximum length of tokens is $1500$. Our decoder-only Transformer consists of $24$ layers with position embedding dimensionality of $d_{pos}=512$, parameter embedding dimensionality of $d_{para}=512$, value embedding dimensionality of $d_{val}=512$, and text feature embedding dimensionality of $d_{f}=512$. We train our model using Adam optimizer, with a learning rate of $10^{-4}$ and a batch size of $4$. The model is trained on a single A6000 GPU for 30 hours.
4 Customized Garment Generation
Figure 4: Examples of pattern completion. Given an incomplete pattern, our method infers reasonable sewing pattern completions with various text prompts.
@@ 4 Customized Garment Generation
4.1 PBR Texture Generation
In some production software commonly utilized by fashion designers, designers often create the texture of the garments after completing the pattern design. For garments, designers usually employ tile-based and physically-based textures such as diffuse, roughness, and normal maps to enhance the realistic appearance of the fabric. Therefore, to generate customized garments, we tailor a pre-trained Stable Diffusion [45] and employ a progressive training approach to generate PBR textures guided by text.
@@ 4.1 PBR Texture Generation
Figure 5: Examples of texture editing. By manually drawing or creating sketches with text prompts, our method facilitates user-friendly texture editing.
Figure 6: Qualitative comparisons on sewing patterns. Major errors on panels are marked with red edges. The inputs of each method from left to right are meshes, images, and texts, respectively, which are shown along with the results.
4.2 Customized Generation through User-friendly Interaction
@@
Following the implementation of generating sewing patterns and textures through text prompts, our framework in practical scenarios enables designers to interact with the generator using natural language instead of relying on dataset-like formatted prompts. We adopt GPT-4 [1] with content learning to interpret users’ natural language inputs, subsequently producing shape prompts and texture prompts. These prompts are then fed to the SewingGPT and the PBR texture generator, respectively. Once sewing patterns are generated, we stitch them onto a T-pose human model. Subsequently, the generated garments, along with PBR textures, seamlessly integrate into industrial software, allowing for animation with the human model and rendering under various lighting, ensuring vivid, realistic results.
@@
Production settings usually necessitate generating multiple clothing items (e.g., daily outfits like pants, T-shirts, and jackets) simultaneously. Past 3D content generation studies based on mesh or implicit fields face challenges in effectively achieving layered draping of multiple garments on a target human model. The adoption of sewing pattern representation enables the respective generation of multiple garments and their natural draping onto the human model.
@@
In the majority of recent 3D generation tasks, the inability to produce structured UV maps has been a significant impediment, particularly for generating garments. However, our generation method, utilizing sewing pattern representation, enables the creation of distinct and structured UV mappings of each panel. This facilitates convenient texture editing at specific locations, allowing efficient post-processing on the textures. As shown in Figure 5, we demonstrate the ShanghaiTech icon drawn on a cream-colored T-shirt’s diffuse map and a duck seamlessly blended with the original grey-colored pants’ diffuse map by creating hand-made sketches with text prompts in the masked areas (refer to stable-diffusion-webui [3] for detailed implementation).
5 Experiments
Figure 7: Our results gallery of DressCode. We generate sewing patterns, PBR textures, and garments in diverse poses and lighting conditions, guided by various text prompts.
In this section, we first conduct qualitative and quantitative comparison experiments with other state-of-the-art 3D generation methods to demonstrate the generation capability of our method. We then present ablation studies to evaluate our pipeline. Furthermore, we conduct a comprehensive user study to showcase our results compared to other methods. Finally, we show a results gallery in Figure 7 of generated high-quality garments using our method with various text prompts, including generated sewing patterns, PBR textures, draping results on a T-pose human model, and animated results with various poses under different illuminations.
Figure 8: Qualitative comparisons on text-to-3D generation. Images on the right side of our methods are the draped garments on human models. Images on the right top of other methods are the top view of each garment. We show our results yield high-quality rendering and the capability to drape on human bodies.
5.1 Comparison of 3D Garment Generation
@@
We show some qualitative comparisons with two state-of-the-art sewing pattern works [22, 26] in Figure 6, where we present the panel prediction, draped garment on a T-pose human model, and corresponding inputs for each method. Given that NeuralTailor is designed for 3D point cloud inputs and trained on open-surface meshes, we utilize a 3D generation method [66] using UDF representation to create meshes conditioned on specific garment categories as inputs (denoted as NeuralTailor*). For Sewformer, designed for image inputs, we utilize DALLE-3 [7] to synthesize input images from text prompts (denoted as Sewformer*). While Sewformer is trained on images of human models wearing both upper and lower garments, we synthesize the images with the same rule and extract partial target panels from predicted results for comparison. Since the generated input meshes are mostly out of the domain of NeuralTailor’s training dataset, the results appear as distorted panels and fail to be stitched together. Sewformer is trained on a new dataset with better generalization; nevertheless, it also encounters issues with irregular and distorted panels, as well as poor garments after stitching. Our method, yielding more accurate results, demonstrates robust generation capabilities with text prompts.
Comparisons on text-to-3D generation.
We evaluate the quality of our customized garment generation with various 3D generation methods in Figure 8. We present qualitative comparisons with two state-of-the-art 3D content generation methods: Wonder3D [30], a 3D creation method from single images, and Richdreamer [41], a text-to-3D work, generating with PBR textures. We adopt DALLE-3 [7] to synthesize image inputs for Wonder3D (denoted as Wonder3D*). Garments generated by Wonder3D fail to retain fine detail and fidelity to the input images, yielding poor geometry. Richdreamer takes approximately 4 hours to yield more realistic results; however, the generated garments are still blurry for rendering. Furthermore, these generated garments are close-surface meshes, as shown in Figure 8, and fail to adapt to human bodies. In contrast, our method facilitates draping garments on human models in various poses and generating high-quality tile-based PBR textures, achieving realistic rendering.
Additionally, we adopt the CLIP score to quantitatively measure different methods. We generate 15 garments with highly diverse text prompts using each method. Then, we render the generated 3D garments with textures and calculate the CLIP score using the given text prompts. The results are shown in Table 1. Although our method doesn’t optimize the 3D models for fitting the rendered images to text prompts better, our model achieves the highest CLIP score, demonstrating the effectiveness of our method. We also compare several characteristics among different methods, highlighting the advantages of our CG-friendly asset generation.
@@
Table 1: Quantitative and characteristic comparisons on different methods. Compared to other methods, our method achieves the highest CLIP score and yields several CG-friendly characteristics.
Figure 9: Ablation on embeddings. Better sewing patterns are generated with our designed parameter embedding and positional embedding.
5.2 Ablation study
@@
5.3 User Study
We conduct a comprehensive user study to evaluate the quality of our generated customized garments, particularly their alignment with given text prompts and overall quality. Given 20 text prompts, including descriptions of the shape and texture of garments, we then render the generated results for each method. Then we ask users to select the best results from all candidates with comprehensive consideration for two aspects: conformity to the prompt texts regarding garments’ shape and texture, and the visualized quality and fidelity of the rendered garments. As illustrated in Table 2, the preference results indicate a significant advantage of our method over competing approaches.
@@ 5.3 User Study
Table 2: Quantitative results of user study. It demonstrates the higher user preference for our method compared to other methods.
6 Limitations and Conclusions
In conclusion, this paper introduces DressCode, a novel customized garment generation framework featuring a sewing pattern generator SewingGPT. DressCode empowers users to effectively create high-quality and CG-friendly garments with sewing patterns and physically-based textures. In contrast to past work, our approach democratizes fashion design through its accessibility and interactivity and enhances digital garments’ practical utility in CG pipelines for post-editing, animation, and realistic rendering. Our experimental results and user studies demonstrate the effectiveness of our method. We envisage our method to benefit CG production and advance the digital garments landscape for virtual try-on, fashion design, and digital human creations.
Despite producing high-quality garment generation, our method encounters certain limitations. The current sewing pattern dataset limits the generation of multi-layered garments (e.g., hoods with pockets), underscoring the importance of dataset expansion to include more complex stitching relationships. Furthermore, the text-driven generation method is subject to biases inherent in underlying pre-trained models such as CLIP and Stable Diffusion. Its user-friendliness and high-quality outputs also carry potential risks for misuse, emphasizing the necessity for future initiatives to tackle these ethical concerns through bias mitigation and thorough review.

### v2

@@ 3.3. Implementation Details
We utilize the extensive sewing pattern garment dataset from (Korosteleva and Lee, 2021), notable for its comprehensive range of sewing patterns and styles of garments, including shirts, hoods, jackets, dresses, pants, skirts, jumpsuits, vests, etc. Our experiments use approximately 19264 samples across 11 fundamental categories. Each garment in the dataset contains a sewing pattern file, a 3D garment mesh draped on a T-pose human model, and a rendered image. We employ the GPT-4V (Achiam et al., 2023) to generate captions for garments from the rendered images of the front view and the back view. For each garment, we first prompt GPT-4V to generate its common name (e.g., hood, T-shirt, blouse) if available, followed by a request for specific geometric features (e.g., long sleeves, wide garment, deep collar), as demonstrated in Figure 4. We combine these two descriptions to form the caption for each garment. In our experiments, we use the pre-defined order of panels in the dataset for training. Additionally, We utilize about 90% of the data from each category for training and the remaining for validation.
@@
We set $K=14$, and the maximum length of tokens is $1500$. Our decoder-only Transformer consists of $24$ layers with position embedding dimensionality of $d_{pos}=512$, parameter embedding dimensionality of $d_{para}=512$, value embedding dimensionality of $d_{val}=512$, and text feature embedding dimensionality of $d_{f}=512$. We set constants $C_{E}=50,C_{R}=1000,C_{T}=1000,C_{S}=1000$ and, $C=1000$. Our CLIP embeddings have a dimension of $d_{\text{CLIP}}=1024$, and condensed feature embeddings have a dimension of $d_{\text{feature}}=512$. We train our model using Adam optimizer, with a learning rate of $10^{-4}$ and a batch size of $4$. The model is trained on a single A6000 GPU for 30 hours.
4. Customized Garment Generation
@@
4.1. PBR Texture Generation
In some production software commonly utilized by fashion designers, designers often create the texture of the garments after completing the pattern design. For garments, designers usually employ tile-based and physically-based textures such as diffuse, roughness, and normal maps to enhance the realistic appearance of the fabric. Therefore, to generate customized garments, we tailor a pre-trained Stable Diffusion (Rombach et al., 2022) and employ a progressive training approach to generate PBR textures guided by text.
@@ 4.1. PBR Texture Generation
4.2. Customized Generation through User-friendly Interaction
@@
Following the implementation of generating sewing patterns and textures through text prompts, our framework in practical scenarios enables designers to interact with the generator using natural language instead of relying on dataset-like formatted prompts. We adopt GPT-4 (Achiam et al., 2023) with content learning to interpret users’ natural language inputs, subsequently producing shape prompts and texture prompts. These prompts are then fed to the SewingGPT and the PBR texture generator, respectively. Once sewing patterns are generated, we stitch them onto a T-pose human model. Subsequently, the generated garments, along with PBR textures, seamlessly integrate into industrial software, allowing for animation with the human model and rendering under various lighting, ensuring vivid, realistic results.
Figure 6. Examples of our multiple garments draping process. Starting with initial sewing patterns, we first drape the inside T-shirt, followed by draping the outside jacket onto the model’s body.
@@
Production settings usually necessitate generating multiple clothing items (e.g., daily outfits like pants, T-shirts, and jackets) simultaneously. Past 3D content generation studies based on mesh or implicit fields face challenges in effectively achieving layered draping of multiple garments on a target human model. The adoption of sewing pattern representation enables the respective generation of multiple garments and their natural draping onto the human model. In our work, for T-pose results, we use the Qualoth simulator (Choi and Ko, 2002) as the physics simulator. We utilize the same material parameters and 3D human model from (Korosteleva and Lee, 2021). In the process of draping multiple garments, we employ an automated sequential multi-garment draping technique, as depicted in Figure 6. Specifically, for a set of clothes, we drape the garment onto the model’s body from the inside out. After each simulation, we combine the mesh of the simulated garment and the human model, then perform the next simulation with the subsequent garment.
Figure 7. Examples of pattern completion. Given an incomplete pattern, our method infers reasonable sewing pattern completions with various text prompts.
@@
Figure 8. Examples of texture editing. By manually drawing or creating sketches with text prompts, our method facilitates user-friendly texture editing.
Figure 9. Our results gallery of DressCode. We generate sewing patterns, PBR textures, and garments in diverse poses and lighting conditions, guided by various text prompts.
Figure 10. Qualitative comparisons on sewing patterns. Major errors on panels are marked with red edges. The inputs of each method from left to right are meshes, images, and texts, respectively, which are shown along with the results.
@@
In the majority of recent 3D generation tasks, the inability to produce structured UV maps has been a significant impediment, particularly for generating garments. However, our generation method, utilizing sewing pattern representation, enables the creation of distinct and structured UV mappings of each panel. This facilitates convenient texture editing at specific locations, allowing efficient post-processing on the textures. As shown in Figure 8, we demonstrate the SIGGRAPH icon drawn on a cream-colored T-shirt’s diffuse map and a duck seamlessly blended with the original grey-colored pants’ diffuse map by creating hand-made sketches with text prompts in the masked areas (refer to stable-diffusion-webui (AUTOMATIC1111, 2022) for detailed implementation).
5. Experiments
In this section, we first conduct qualitative and quantitative comparison experiments with other state-of-the-art 3D generation methods to demonstrate the generation capability of our method. We then present ablation studies to evaluate our pipeline. Furthermore, we conduct a comprehensive user study to showcase our results compared to other methods. We also verify our results in Marvelous Designer (CLO3D, 2024); we manually load generated garments into the software and then animate the garments with the human model in non-T-pose. Finally, we show a results gallery in Figure 9 of generated high-quality garments using our method with various text prompts, including generated sewing patterns, PBR textures, draping results on a T-pose human model, and animated results with various poses under different illuminations.
Figure 11. Qualitative comparisons on text-to-3D generation. Images on the right side of our methods are the draped garments on human models. Images on the right top of other methods are the top view of each garment. We show our results yield high-quality rendering and the capability to drape on human bodies.
5.1. Comparison of 3D Garment Generation
@@
We show some qualitative comparisons with two state-of-the-art sewing pattern works (Korosteleva and Lee, 2022; Liu et al., 2023d) in Figure 10, where we present the panel prediction, draped garment on a T-pose human model, and corresponding inputs for each method. Given that NeuralTailor is designed for 3D point cloud inputs and trained on open-surface meshes, we utilize a 3D generation method Surf-D (Yu et al., 2023) using the UDF representation to create meshes conditioned on specific garment categories as inputs (denoted as NeuralTailor*). Note that Surf-D is trained on the Deep Fashion3D dataset. Although our method supports complex prompts, we only select category names such as skirt and sleeveless dress, which are presented in the Deep Fashion3D dataset, in our experiment as prompts to ensure a fair comparison. For Sewformer, designed for image inputs, we utilize DALLE-3 (Betker et al., 2023) to synthesize input images from text prompts (denoted as Sewformer*). While Sewformer is trained on images of human models wearing both upper and lower garments, we synthesize the images with the same rule and extract partial target panels from predicted results for comparison. For the first row of Figure 10, our generated image involves a model wearing both a top shirt and a skirt, not only a sole skirt, which aligns with the training dataset in Sewformer. For fairness, we manually extract skirt patterns from Sewformer predictions, as shown in the comparison. Since the generated input meshes are mostly out of the domain of NeuralTailor’s training dataset, the results appear as distorted panels and fail to be stitched together. Sewformer is trained on a new dataset with better generalization; nevertheless, it also encounters issues with irregular and distorted panels, as well as poor garments after stitching. Our method, yielding more accurate results, demonstrates robust generation capabilities with text prompts.
@@
Table 1. Quantitative and characteristic comparisons on different methods. Compared to other methods, our method achieves the highest CLIP score and yields several CG-friendly characteristics.
Comparisons on text-to-3D generation.
We evaluate the quality of our customized garment generation with various 3D generation methods in Figure 11. We present qualitative comparisons with two state-of-the-art 3D content generation methods: Wonder3D (Long et al., 2023), a 3D creation method from single images, and RichDreamer (Qiu et al., 2023), a text-to-3D work, generating with PBR textures. We adopt DALLE-3 (Betker et al., 2023) to synthesize image inputs for Wonder3D (denoted as Wonder3D*). Wonder3D takes about 4 minutes to generate garments but fails to retain fine detail and fidelity to the input images, yielding poor geometry. RichDreamer takes approximately 4 hours to optimize and yield more realistic results; however, the generated garments are still blurry for rendering. Furthermore, these generated garments are close-surface meshes, as shown in Figure 11, and fail to adapt to human bodies. In contrast, our method takes about 1 minute to generate sewing patterns, and overall about 3 minutes to generate the simulated garments. It facilitates draping garments on human models in various poses and generating high-quality tile-based PBR textures, achieving realistic rendering.
Additionally, we adopt the CLIP score to quantitatively measure different methods. We generate 15 garments with highly diverse text prompts using each method. Then, we render the generated 3D garments with textures and calculate the CLIP score using the given text prompts. Although our method does not optimize the 3D models for fitting the rendered images to text prompts better, our model achieves the highest CLIP score, demonstrating the effectiveness of our method. These general 3D methods are more broadly applicable than to only 3D garments. However, we also compare several characteristics among different methods, highlighting the advantages of our CG-friendly asset generation in the specific 3D garment domain. The results are shown in Table 1.
Figure 12. Ablation on embeddings. Better sewing patterns are generated with our designed parameter embedding and positional embedding.
Figure 13. Validation. We validate our generated garments with the text prompts from our dataset out of training and in the wild. Our results align well with the text prompts, demonstrating the effectiveness of our method.
5.2. Ablation Study
@@
5.3. Validation
To validate our methods, we design two experiments, as shown in Figure 13. Firstly, we test text prompts from the holdout set as done for training and compare the results in the dataset with our generated outputs. In the first two rows of Figure 13, we test two data points from outside the training dataset and observe that the generated results correspond well to the input prompts. Notably, in the dress example in the first row, although our generated result includes a waistband while the reference example does not, this still aligns well with the “fitted waist” attribute mentioned in the prompt. Furthermore, we test an image in the wild from the Deep Fashion3D dataset (Zhu et al., 2020) and utilize the method described in Section 3.3 with GPT-4V (Achiam et al., 2023) to generate the caption. This serves as the input to qualitatively compare our result with the reference image, revealing that our result closely resembles the input image. This demonstrates that our work effectively bridges the gap between conceptual textual descriptions and their practical, visual counterparts in garment design.
Figure 14. Qualitative comparison with Parametric Templates. The user inputs the prompt “a loose night-robe,” and we show the results from Parametric Templates and our proposed SewingGPT.
5.4. Comparison with Parametric Templates
(Korosteleva and Lee, 2021) propose a method of a flexible description structure for specifying parametric sewing pattern templates. An intuitive idea is that we can control the parameters of the pre-defined parametric sewing pattern templates to generate diverse garments (denoted as Parametric Templates). To facilitate interaction through natural language and benefit from the strong capability of content learning in ChatGPT, we employ GPT-4 (Achiam et al., 2023) in our experiment, designing prompts to enable its role as a garment design assistant, providing formatted outputs. Subsequently, we inquire about the description of garments, prompting it to output parameters for a specific template. As illustrated in the middle left of Figure 14, GPT-4 responds with several parameters when we ask with “a loose night-robe.” We qualitatively compare the results from Parametric Templates and our SewingGPT in Figure 14, observing that both methods generate reasonable results. Nevertheless, our proposed SewingGPT is more adaptable to the diverse categories of data, as it does not require selecting pre-defined templates and designing prompts specifically for ChatGPT’s content learning. Additionally, our methods enable us to extend beyond such parametrized datasets for more complex sewing patterns.
Figure 15. Examples from our user study. We present two cases for users to select the best results from three options: A, B, and C. The left group of results is generated by Ours, RichDreamer, and Wonder3D*, respectively, and the right group of results is generated by Wonder3D*, Ours, and RichDreamer, respectively. The generation method of each image is not disclosed to the users.
@@ 5.4. Comparison with Parametric Templates
Table 2. Quantitative results of user study. It demonstrates the higher user preference for our method compared to other methods.
5.5. User Study
We conduct a comprehensive user study to evaluate the quality of our generated customized garments, particularly their alignment with given text prompts and overall quality. Given 20 text prompts, including descriptions of the shape and texture of garments, we then render the generated results for each method and shuffle the order of results obtained by different methods, as shown in Figure 15. Then we ask approximately 30 users to select the best results from all candidates with comprehensive consideration for two aspects: conformity to the captions of prompt texts regarding garments’ shape and texture, and the visualized quality and fidelity of the rendered garments. As illustrated in Table 2, the preference results clearly indicate a significant advantage of our method over competing approaches in both aspects, highlighting its superiority in aligning closely with the textual prompts and producing visually appealing and high-fidelity garments.
5.6. Limitations and Discussions
Despite producing high-quality garment generation from text guidance, our method encounters certain limitations. One limitation is that the current sewing pattern dataset limits the generation of multi-layered garments, such as “hoodie jacket with a pocket,” as depicted in the first row of Figure 16. It underscores the importance of dataset expansion to include more complex stitching relationships. Another limitation is that our model struggles with prompts outside the domain of our dataset. For instance, we test prompts like “one-shoulder dress.” As shown in the second row of Figure 16, the model still generates a “two-shoulder dress,” due to the absence of “one-shoulder” garments in our dataset, which hinders its ability to recognize this attribute. Similarly, we experiment with integrating unusual characteristics into garments, combining specific attributes from different categories of garments, such as a “dress with a hood,” a style not commonly encountered in real life. Our results shown in the last row of Figure 16 display a very loose hoodie jacket. Although the results somewhat resemble a dress with its loose style, a dress should not be open-front. This outcome is due to the presence of only hoodie jackets as hooded garments in our dataset, leading to a bias toward producing results within the hoodie jacket category when the prompt includes a “hood.” We believe that enriching the dataset with a wider variety of garments can significantly enhance the model’s versatility. Additionally, inspired by recent breakthroughs in the generation domain, distilling knowledge from the pre-trained foundation model, such as SDS (Poole et al., 2022), to improve generalization is a worthy direction for future work. Lastly, although our framework is pioneering in generating garments with text prompts, incorporating multi-modality inputs could prove more effective. Generating sewing patterns and textures controlled by both text and images presents a particularly intriguing and challenging problem yet to be addressed in real-world applications.
Figure 16. Failure cases. Our method struggles to generate garments that fall outside the domain of the training dataset. We present three examples of generated results: a “hoodie jacket with a pocket,” a “one-shoulder dress,” and a “dress with a hood.” The reference images are generated by DALLE-3 (Betker et al., 2023).
Potential ethical implications.
The text-driven generation method is subject to biases inherent in underlying pre-trained models such as CLIP and Stable Diffusion. Its user-friendliness and high-quality outputs also carry potential risks for misuse, emphasizing the necessity for future initiatives to tackle these ethical concerns through bias mitigation and thorough review. Additionally, the utilization of Stable Diffusion for both fine-tuning and inference purposes raises significant concerns regarding potential copyright issues, as the model may inadvertently generate content that mirrors proprietary works without explicit authorization. It is important for future work to ensure that the training data and generated content of these models are carefully reviewed and selected.
6. Conclusions
In conclusion, this paper introduces DressCode, a novel customized garment generation framework featuring a text-driven sewing pattern generator SewingGPT. This framework democratizes garment design by making it accessible and interactive, enabling both novices and experts to generate detailed sewing patterns and high-quality PBR textures through simple textual prompts. Additionally, our framework supports interaction-friendly applications for garment generation, completion, and editing, providing powerful tools that enable designers to unleash their creative imagination. Our experimental results and user study demonstrate the effectiveness of our method in producing CG-friendly garments that excel in quality and alignment with input prompts.
In contrast to past work, our approach democratizes fashion design through enhanced accessibility and interactivity and improves the practical utility of digital garments in CG pipelines for post-editing, animation, and realistic rendering. The ease of use and innovative approach of DressCode promise exciting developments for future advancements in digital garments, potentially transforming digital garment creation and customization. We envision our method benefiting CG production and advancing the digital garment landscape for virtual try-on, fashion design, and digital human creation.
Acknowledgements.
This work was supported by the National Key R&D Program of China (2022YFF0902301), NSFC programs (61976138, 61977047), STCSM (2015F0203-000-06), and SHMEC(2019-01-07-00-01-E00003). We also acknowledge support from the Shanghai Frontiers Science Center of Human-centered Artificial Intelligence (ShangHAI).